OpenStudy (anonymous):

What are the solutions of 2x2 + 8x = -26? A) -2 ± 3 i B) -2 ± i C) 2 ± 6 i D) -3 ± 2 i

OpenStudy (anonymous):

1) divide all by \(2\) what do you get ?

OpenStudy (anonymous):

i mean starting \[2x^2 + 8x = -26\]divide all by \(2\)

OpenStudy (anonymous):

x1+x2=-b/a

OpenStudy (anonymous):

x1.x2=c/a

OpenStudy (anonymous):

ax²+bx+c

OpenStudy (anonymous):

x^2 + 4x = -13 @satellite73

OpenStudy (anonymous):

k good only two more steps what is half of \(4\) ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

x1+x2=-4 so only can be A or B

OpenStudy (anonymous):

right we go right from \[x^2+4x=-13\] to \[(x+2)^2=-13+2^2\] or \[(x+2)^2=-9\]

OpenStudy (anonymous):

9+4=13 so its A)

OpenStudy (anonymous):

then \[x+2=\pm\sqrt{-9}\\ x+2=\pm3i\\ x=-2\pm3i\]

OpenStudy (anonymous):

so it would be A (: @satellite73 @RaphaelFilgueiras

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you both so so much

OpenStudy (anonymous):

yw from us both

OpenStudy (anonymous):

yw